OpenStudy (anonymous):

How long will a $9000 investment take to earn $180 interest at an annual interest rate of 8%?

OpenStudy (lstrasmann):

I = PRT 180 = (9000)(.08)(x) 180 = 720x 180 + -720x = 720x + -720x 180 + -720x = 0 180 + -180 + -720x = -180 -720x = -180 x = 0.25 (4 months)